An airplane is in level flight at a velocity of 300 miles per hour and an altitude of 5000ft when I drops a bomb. (a) how long does the bomb take to reach the ground? (b) what will the bomb's velocity be on impact?


1
Expert's answer
2021-10-15T10:22:44-0400

(a) The time is


"t=\sqrt{2h/g}=\sqrt{2·5000/32}=17.7\text{ s}."

(b) The horizontal component of velocity does not change (if we ignore air resistance) and it is 440 ft/s. The vertical component:


"v_y=\sqrt{2gh}=566\text{ ft/s},\\\space\\
v=\sqrt{v_x^2+v_y^2}=717\text{ ft/s, or 488 mph.}"